OpenStudy (anonymous):

Lydia received a parking ticket because she parked 10 feet from a fire hydrant. She had estimated her distance from the fire hydrant as being the minimum allowable distance, which is 15 feet. What was her percent error? a] 15% b] 33% c] 50% d] 67%

OpenStudy (a_clan):

b

OpenStudy (anonymous):

The percentage will be (15 - 10)/15 = 1/3 = 33%

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

{[(experimental value)-(theoretical value)]/(theoretical value)}*100 [(15ft-10ft)/(10ft)]*100=50%

OpenStudy (anonymous):

which one would be the theoretical value and which the experimental? i figure since the 15ft was a result of her action it is experimental, and since the required minimum space is 10ft that would be the theoretical value.

OpenStudy (anonymous):

nevermind lol

OpenStudy (a_clan):

15 should be theoretical. It is the allowable distance

OpenStudy (anonymous):

I don't think that's correct, jmw417. |(estimate - actual)/actual| is the percentage error. The error you make must be related to the actual value of the variable,

OpenStudy (anonymous):

yes i fell victim to speed reading on this one. i see that the 10 ft is her experimental value and the 15 ft is the theoretical allowable distance. My bad folks!

OpenStudy (anonymous):

That's ok. No harm done, I am sure :)

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

It isn't 33%. I tried that and it marked me wrong.

OpenStudy (anonymous):

the answer would be 50 %